OpenStudy (anonymous):

Solve the following equation. -3 = z + 4 I will give a medal! Please help and thank you!

OpenStudy (anonymous):

-4 - 3 = Z -7 = Z

OpenStudy (anonymous):

You will first subtract four to both sides of the equation And your answer is z = -7

OpenStudy (anonymous):

Thank you!